40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14078     Entry time: Tue Jul 17 17:37:46 2018
Author: Annalisa, Terra 
Type: Configuration 
Category: Thermal Compensation 
Subject: Heaters installation 

Summary

We installed two heaters setup on the ETMY bench in order to try inducing some radius of curvature change and therefore HOMs frequency shift.

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

We installed two heaters setup.

Elliptic reflector setup (H1): heater put in the focus of the elliptical reflector: this will make a heat pattern as descirbed in the elogs #14043 and #14050. 

 

Lenses setup (H2): heater put in a cylndrical reflector (made up with aluminum foil) 1'' diameter, and 2 ZnSe lenses telescope, composed by a 1.5'' and a 1'' diameter respectively, both 3.5'' focal length. The telescope is designed in such a way to focus the heat map on the mirror HR surface. For this latter the schematic was supposed to be the following: 

This setup will project on the mirror a heat pattern like this:

which is very convenient if we want to see a different radius of curvature for different HOMs. However, the power that we are supposed to have absorbed by the mirror with this setup is very low (order of 40-ish mW with 18V, 1.2A) which is probably not enough to see an effect. Moreover, mostly for space reasons (post base too big), the distances were not fully kept, and we ended up with the following setup:

In this configuration we won't probably have a perfect focusing of the heat pattern on the mirror.

In vacuum connections

See Koji's elog #14077 for the final pin connection details. In summary, in vacuum the pins are:

13 to 8 --> cable bunch 0

7 to 2 --> cable bunch 2

25 to 20 --> cable bunch 1

19 to 14 --> cable bunch 3

where Elliptic reflector setup (H1) is connected to cables 0 and 1, and the lenses setup is connected to cables 2 and 3.

Installed setup

This is the installed setup as seen from above:

Attachment 5: IMG_5380.JPG  4.954 MB  Uploaded Wed Jul 18 23:21:44 2018  | Show | Hide all | Show all
ELOG V3.1.3-